# Many Heads Are Better Than One: Improved Scientific Idea Generation by A LLM-Based Multi-Agent System

### v2

@@ 4.1 Experimental Settings
Datasets. We evaluate the performance of VirSci and baseline methods on two datasets: the AMiner Computer Science dataset [^5] https://www.aminer.cn/aminernetwork, containing 1,712,433 authors and 2,092,356 papers from 1948 to 2014, and the Open Academic Graph 3.1666https://open.aminer.cn/open/article?id=65bf053091c938e5025a31e2, comprising 35,774,510 authors and 130,710,733 papers as of 2023. For quality assurance, we applied filtering strategies, reducing the Computer Science dataset to 156 authors and 178,197 papers, and the Open Academic Graph to 3,169 authors and 201,131 papers. Additional dataset details and filtering strategies are provided in Appx. B.
Evaluation Metrics.  In this paper, we adopt a more objective approach by using three common metrics that align with our intuition, as no single metric fully captures the novelty of scientific outputs: (1) Historical Dissimilarity (HD): The average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in $B_{\text{past}}$ (shao2020bert, ; zhou2024fine, ). A larger distance indicates greater dissimilarity from existing papers, suggesting a higher likelihood of novelty. (2) Contemporary Dissimilarity (CD): The average Euclidean distance between the generated abstract embedding and embeddings of the top 5 most similar abstracts in $B_{\text{con}}$. A smaller distance indicates greater similarity to newer papers, also suggesting a higher likelihood of novelty. (3) Contemporary Impact (CI): The average citation count of the top 5 most similar abstracts in $B_{\text{con}}$ (yang2022gender, ). A higher citation count suggests that the generated abstract is more likely to have a higher impact. To ensure comparability, we normalize each calculated metric using the mean value taken over all papers published in the same year as the chosen similar abstract in the corresponding database, with normalization defined as the metric divided by the mean value (alshebli2018preeminence, ).
@@
Baselines. In this section, we compare VirSci with two recent agent-based systems for scientific discovery: the multi-agent system HypoGen qi2024large and the SOTA single-agent system AI Scientist lu2024ai . Given the differences in their pipelines, we adjust the experimental settings to a comparable level for a fair evaluation, elaborated in Appx. B.3. We then assess the results using our proposed metrics, LLM review scores, and human evaluations (Appx. B.4) to demonstrate the superiority of our method in idea generation.
4.2 Analysis of Proposed Metrics
To assess the validity of our proposed overall novelty (ON) metric, we selected 200 abstracts generated from the Computer Science dataset, which were evaluated using both ON and by human experts in the field of computer science (details in Appx. B.4). For the human evaluations, we the novelty scoring framework from (si2024can, ) as a guideline. To ensure objectivity, independent researchers assessed each abstract, and the average score was calculated. As shown in Fig. 3, the Pearson correlation coefficient between ON and human ratings demonstrates a positive correlation, which supports the validity of our metric. We also explore the correlation between ON and LLM-based review (a common novelty measurement method (gu2024survey, )) in Appx. C and conduct further analysis.
Method LLM $\uparrow$ Proposed Metric Human Evaluation
CD $\downarrow$ CI $\uparrow$ Nov $\uparrow$ Fea $\uparrow$ Eff $\uparrow$
@@ 4.2 Analysis of Proposed Metrics
4.3 Comparisons with Baselines
4.4 Exploring Collaboration Mechanism
While the dynamics of group collaboration have been extensively studied in the Science of Science, its applicability to artificial multi-agent systems and the effects it may produce remain unclear. In this section, we analyze factors influencing idea novelty in our system, including team size, freshness, and research diversity—factors previously analyzed in human teams (wu2019large, ; zeng2021fresh, ; shi2023surprising, ).
Effects of Team Size on Novelty. The results in Fig. 4 show that increasing team size can enhance the ON of generated abstracts by bringing in diverse ideas and perspectives. However, this relationship is not strictly linear, with peak ON occurring at a team size of 8. While moderate team expansion boosts novelty, excessively large teams can create coordination and communication challenges, leading to diluted contributions and groupthink. This aligns with existing literature, which suggests that smaller teams generate more disruptive ideas, while larger teams focus on refining existing concepts(wuchty2007increasing, ; fortunato2018science, ; wu2019large, ).
Effects of Discussion Turn on Novelty.  The number of discussion turns plays a crucial role in fostering novel ideas (mezirow2003critical, ; li2023camel, ; shinn2024reflexion, ; lu2024ai, ). As shown in our experimental results (Fig. 4), an optimal number of turns promotes deeper insights and more innovative research outputs. While the initial turns facilitate idea generation, excessive turns can result in fatigue and reduced engagement, ultimately hindering creativity. Our findings suggest that peak ON is achieved at around 5 discussion turns.
Consistency Between Two Datasets. As shown in Fig. 4, no significant differences were found in outputs from single-agent teams (team size equals 1) across the two datasets, as individual research lacks interdisciplinary input. The overall trend—novelty rising, peaking at a team size of 8 and 5 turns, then declining—was consistent across both datasets, demonstrating our platform’s robustness and the potential of multi-agent systems in enhancing scientific idea generation. These findings also support the value of interdisciplinary collaborations for driving higher-impact research (shi2023surprising, ).
@@ 4.4 Exploring Collaboration Mechanism
Effects of Team Freshness on Novelty. As shown in Fig. 5, team freshness—the fraction of members who have not previously collaborated—affects the novelty of outputs. Freshness has its strongest effect at 50%, especially for larger teams (size 8), where HD peaks. This suggests that a balanced mix of new and returning members promotes innovation by diverging from past research. As freshness increases, CD decreases, indicating that teams with more fresh members align their ideas more closely with future research trends. Both CI and ON reach their highest values around 50% freshness, highlighting that a balanced team optimally combines novelty and future relevance, consistent with previous work (guimera2005team, ; zeng2021fresh, ).
@@
Effects of Team Research Diversity on Novelty. Team research diversity is defined as the proportion of team members specializing in distinct research topics. As shown in Fig. 6, moderate diversity boosts HD, with peak performance at 25-50% diversity for 4-member teams and 50-75% for 8-member teams across both datasets. CD drops before reaching 50% diversity, suggesting that diverse teams align better with emerging research trends while maintaining innovation. Larger teams benefit more from higher diversity, showing an increase in CI, while smaller teams exhibit more stable, moderate effects. Overall, ON follows a reverse U-shaped curve for both team sizes and datasets, underscoring the importance of balancing research diversity for novel outcomes. This conclusion mirrors findings in Science of Science, where unexpected team combinations can enhance research impact (uzzi2013atypical, ; shi2023surprising, ).
5 Ablation Study
5.1 Effects of Paper Database
@@
In the workflow of VirSci, the past paper database is utilized as a reference during idea generation and novelty assessment. To verify the role of the constructed paper database, we conduct ablation studies on both datasets under the setting of 8 team members and 5 discussion turns, as shown in Tab. 2 and 3. The experimental results demonstrate that having references enhances the novelty of the final generated ideas, avoiding shallow or fanciful outcomes produced by the system.
5.2 Effects of Components Designed for Improving Novelty
Team Size Invitation Turns
@@ 5.2 Effects of Components Designed for Improving Novelty
Team Size Assessment Turns
@@ 5.2.2 Novelty Assessment Mechanism
5.2.3 Self-review Mechanism
@@
5.3 Effects of Discussion Pattern
5.3.1 Discussion Topologies
Team Size Topology Turns
1 2 3 4 5 6 7
4 Sequential 1.75 0.90 2.36 2.00 3.40 2.91 2.53
Random 1.67 0.85 2.34 2.01 3.35 2.84 2.47
8 Sequential 3.48 3.67 3.97 3.56 4.23 3.48 3.43
Random 3.50 3.63 3.88 3.52 4.16 3.40 3.39
5.3.2 Discussion Turns
[…]
@@ 5.3.2 Discussion Turns
In the previous experiments, we fixed the number of discussion turns in each step to ensure fair comparisons. However, in real-world research environments, teams of scientists do not spend the same amount of time on each stage of the research process. To explore this, we compare fixed discussion turns with adaptive turn numbers. In the adaptive pattern, the team leader decides whether the team needs additional turns based on the current progress and the goals of each stage. The results of both patterns, along with their corresponding inference cost, are shown in Tab. 8.
The comparison reveals that the adaptive pattern achieves a higher ON while reducing the inference cost. This efficiency can be attributed to the more flexible approach, allowing teams to adjust their discussions dynamically rather than adhering to a rigid structure (which may lead the team in the wrong direction when a section is over-discussed). Furthermore, examining the number of turns at each stage in both 4-person and 8-person teams under the adaptive pattern offers additional insights. Larger teams require more discussion turns and face greater challenges in reaching consensus (janis1972victims, ; pitters2014psychology, ). This highlights the adaptive pattern’s advantage in accommodating the complexities of larger teams while maintaining a higher level of novelty in final research outputs.
5.4 Effects of Different Exploration Mechanisms on Scientific Collaboration
If we regard the collaboration among our scientists as an ”explore-exploit” model, collaborating with scientists they have not worked with before can be viewed as “explore”, while collaborating with previously partnered scientists can be seen as “exploit”. In this section, we further investigate how the mechanism of ”explore” impacts our final results.
In the default setting, when creating the adjacency matrix, we add 1 to each value in the matrix to ensure that agents collaborate with individuals they have not worked with before, treating ”explore” as uniformly distributed. Based on this, we also conducted experiments where the incremental values follow a normal distribution, defined as:

### v3

@@ 4.1 Experimental Settings
Datasets. We evaluate the performance of VirSci and baseline methods on two datasets: the AMiner Computer Science dataset [^3] https://www.aminer.cn/aminernetwork, containing 1,712,433 authors and 2,092,356 papers from 1948 to 2014, and the Open Academic Graph 3.1444https://open.aminer.cn/open/article?id=65bf053091c938e5025a31e2, comprising 35,774,510 authors and 130,710,733 papers as of 2023. For quality assurance, we applied filtering strategies, reducing the Computer Science dataset to 156 authors and 178,197 papers, and the Open Academic Graph to 3,169 authors and 201,131 papers. Additional dataset details and filtering strategies are provided in Appx. C.
Evaluation Metrics.  In this paper, we adopt a more objective approach by using three common metrics that align with our intuition, as no single metric fully captures the novelty of scientific outputs: (1) Historical Dissimilarity (HD): The average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in $B_{\text{past}}$ (Shao et al., 2020; Zhou et al., 2024). A larger distance indicates greater dissimilarity from existing papers, suggesting a higher likelihood of novelty. (2) Contemporary Dissimilarity (CD): The average Euclidean distance between the generated abstract embedding and embeddings of the top 5 most similar abstracts in $B_{\text{con}}$. A smaller distance indicates greater similarity to newer papers, also suggesting a higher likelihood of novelty. (3) Contemporary Impact (CI): The average citation count of the top 5 most similar abstracts in $B_{\text{con}}$ (Yang et al., 2022). A higher citation count suggests that the generated abstract is more likely to have a higher impact. To ensure comparability, we normalize each calculated metric using the mean value taken over all papers published in the same year as the chosen similar abstract in the corresponding database, with normalization defined as the metric divided by the mean value (AlShebli et al., 2018).
@@
[…]
@@
Baselines. In this section, we compare VirSci with two recent agent-based systems for scientific discovery: the multi-agent system HypoGen Qi et al. (2024) and the SOTA single-agent system AI Scientist Lu et al. (2024). Given the differences in their pipelines, we adjust the experimental settings to a comparable level for a fair evaluation, elaborated in Appx. C.3. We then assess the results using our proposed metrics, LLM review scores, and human evaluations (Appx. C.4) to demonstrate the superiority of our method in idea generation.
4.2 Analysis of Proposed Metrics
To assess the validity of our proposed overall novelty (ON) metric, we selected 200 abstracts generated from the Computer Science dataset, which were evaluated using both ON and by human experts in the field of computer science (details in Appx. C.4). For the human evaluations, we the novelty scoring framework from (Si et al., 2024) as a guideline. To ensure objectivity, independent researchers assessed each abstract, and the average score was calculated. As shown in Fig. 3, the Pearson correlation coefficient between ON and human ratings demonstrates a positive correlation, which supports the validity of our metric. We also explore the correlation between ON and LLM-based review (a common novelty measurement method (Gu et al., 2024)) in Appx. E and conduct further analysis.
4.3 Comparisons with Baselines
4.4 Exploring Collaboration Mechanism
While the dynamics of group collaboration have been extensively studied in the Science of Science, its applicability to artificial multi-agent systems and the effects it may produce remain unclear. In this section, we analyze factors influencing idea novelty in our system, including team size, freshness, and research diversity—factors previously analyzed in human teams (Wu et al., 2019; Zeng et al., 2021; Shi and Evans, 2023).
@@ 4.4 Exploring Collaboration Mechanism
Effects of Team Size on Novelty. The results in Fig. 4 show that increasing team size can enhance the ON of generated abstracts by bringing in diverse ideas and perspectives. However, this relationship is not strictly linear, with peak ON occurring at a team size of 8. While moderate team expansion boosts novelty, excessively large teams can create coordination and communication challenges, leading to diluted contributions and groupthink. This aligns with existing literature, which suggests that smaller teams generate more disruptive ideas, while larger teams focus on refining existing concepts(Wuchty et al., 2007; Fortunato et al., 2018; Wu et al., 2019).
Effects of Discussion Turn on Novelty.  The number of discussion turns plays a crucial role in fostering novel ideas (Mezirow, 2003; Li et al., 2023; Shinn et al., 2024; Lu et al., 2024). As shown in our experimental results (Fig. 4), an optimal number of turns promotes deeper insights and more innovative research outputs. While the initial turns facilitate idea generation, excessive turns can result in fatigue and reduced engagement, ultimately hindering creativity. Our findings suggest that peak ON is achieved at around 5 discussion turns.
@@
Effects of Team Freshness on Novelty. As shown in Fig. 5, team freshness—the fraction of members who have not previously collaborated—affects the novelty of outputs. Freshness has its strongest effect at 50%, especially for larger teams (size 8), where HD peaks. This suggests that a balanced mix of new and returning members promotes innovation by diverging from past research. As freshness increases, CD decreases, indicating that teams with more fresh members align their ideas more closely with future research trends. Both CI and ON reach their highest values around 50% freshness, highlighting that a balanced team optimally combines novelty and future relevance, consistent with previous work (Guimera et al., 2005; Zeng et al., 2021).
@@
Effects of Team Research Diversity on Novelty. Team research diversity is defined as the proportion of team members specializing in distinct research topics. As shown in Fig. 6, moderate diversity boosts HD, with peak performance at 25-50% diversity for 4-member teams and 50-75% for 8-member teams across both datasets. CD drops before reaching 50% diversity, suggesting that diverse teams align better with emerging research trends while maintaining innovation. Larger teams benefit more from higher diversity, showing an increase in CI, while smaller teams exhibit more stable, moderate effects. Overall, ON follows a reverse U-shaped curve for both team sizes and datasets, underscoring the importance of balancing research diversity for novel outcomes. This conclusion mirrors findings in Science of Science, where unexpected team combinations can enhance research impact (Uzzi et al., 2013; Shi and Evans, 2023).
5 Conclusion
6 Limitations
While our system effectively models scientific collaboration, it simplifies the complexities of real-world teamwork. In large research communities, multiple teams often work on related projects either collaboratively or independently, and researchers frequently participate in multiple teams simultaneously. Our current approach does not fully capture these intricate dynamics, potentially limiting its ability to model real-world scientific collaboration with high fidelity.
Beyond structural limitations, our system also inherits biases from the underlying language model and training data. If not properly mitigated, these biases could disproportionately favor well-established research domains while marginalizing emerging or underrepresented areas. This could reinforce existing disparities in scientific knowledge production, inadvertently shaping the trajectory of research based on historical patterns rather than fostering novel discoveries. Future work should explore methods to detect and counteract these biases, such as dataset diversification and bias-aware fine-tuning Ansar and Talavera (2025).
Another critical limitation is the potential for generating plausible but incorrect or misleading scientific claims. This risk is particularly concerning in high-stakes fields such as medicine and policy-related research, where inaccuracies could lead to misinformation, flawed decision-making, or ethical concerns. Implementing stricter validation mechanisms, such as integrating fact-checking modules or leveraging expert human reviewers, will be necessary to ensure that generated content aligns with established scientific knowledge.
Additionally, our system could be misused for unethical purposes, such as automating the creation of fraudulent research papers or facilitating plagiarism. The ability to generate text that mimics academic writing raises concerns about the proliferation of low-quality or deceptive publications. To mitigate these risks, future work should explore safeguards such as watermarking AI-generated text Zhao et al. (2023), developing robust detection algorithms for synthetic research content, and establishing ethical guidelines for responsible use.
Finally, the biases embedded in large-scale language model training data may disproportionately impact specific research communities. Researchers from underrepresented regions, institutions, or disciplines may find their work underemphasized or misrepresented, reinforcing systemic inequities in academia. Addressing these challenges requires continuous monitoring of dataset composition, proactive inclusion of diverse research sources, and collaboration with domain experts to refine training methodologies.
Our future research will also aim to scale up VirSci to a societal level, incorporating more dynamic and concurrent team interactions. Allowing agents to engage in multiple projects and collaborate across different teams would improve the realism of our simulations and better reflect the complexity of modern scientific collaboration. These advancements would also strengthen the system’s utility for the Science of Science community, enabling deeper investigations into how collaboration influences innovation and knowledge production.
7 Ethics Statement
This research uses two publicly available datasets provided by AMiner, ensuring compliance with data privacy policies. Our system is designed to augment, not replace, human researchers, highlighting the need for human oversight to maintain the quality and integrity of generated outputs. We comply with the licensing terms of the LLMs used, as specified in their official terms of service. To promote research transparency, we have shared all relevant codes for reproducibility within the research community.
While our system supports scientific discovery, it carries potential risks. Data privacy concerns may arise even with public datasets if data handling is improper. To mitigate this, we have applied data anonymization to the datasets, such as masking real scientists’ names in simulations, to prevent data leakage and protect privacy. Additionally, stringent protocols have been implemented to minimize unintended discussions and potential misuse.
Acknowledgement
Appendix A Effect of the Potential Data Leakage
To directly assess the impact of temporal data leakage, we conducted an additional experiment. Using Open Academic Graph 3.2  [^5] https://open.aminer.cn/open/article?id=67aaf63af4cbd12984b6a5f0, we tested a temporal isolation scenario where papers from 2010–2020 were retained as the Past Paper Database, while 30,000 papers from 2024–2025 were used as the new Contemporary Paper Database. Given that LLaMA 3.1’s training data extends only until December 2023, this setup ensured that no "future" papers beyond the model’s training scope were accessible. The experiments were conducted with a team size of 4 and LLaMA3.1-8b, and citation counts were normalized according to the method described in our paper to remove the influence of publication years.
As illustrated in Table 2, when shifting the Contemporary Paper Database from 2021–2023 to 2024–2025, the novelty scores exhibited only a slight decrease, with values remaining consistent across different turns. This suggests that knowledge leakage from future papers does not significantly impact our conclusions.
Appendix B Methodological Details
B.1 Self-review
A self-review mechanism is considered after $R_{abstract}$ is finalized to pre-check its novelty. In this self-review, the optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$. The prompt is:
If the abstract undergoes a second self-review and still does not meet the novelty requirement, it will be discarded, and the team will generate a new idea. Once the self-review yields satisfactory results, the final abstract will be produced, and the system will terminate. However, this self-review mechanism introduces uncertainty in total inference cost, making it difficult to ensure fair experimental comparisons. We discuss the effectiveness of this module only in the ablation study (see Appx. D.2).
Appendix C Experimental Settings
C.1 Implementation
We implement our system on top of the Agentscope framework (Gao et al., 2024), which serves for LLM-empowered multi-agent applications. We evaluate our system using different publicly available LLMs: GPT-4o (OpenAI, 2023), LLaMA3.1-8b (Dubey et al., 2024), and LLaMA3.1-70b. GPT-4o is accessible exclusively via a public API, while the LLaMA3.1 models are open-weight and invoked using the Ollama (Ollama, 2024) in our experiments. LLaMA3.1-8b is chosen as the default LLM considering both efficiency and capability, where more experiments using different LLMs are shown in Appx. D.5. Each experimental run on LLaMA3.1-8b takes approximately 10 minutes on 1 NVIDIA A100 40G GPU within a team discussion setting of 4 members and 5 turns ($K$ = 5). All experimental results are averaged on 20 runs.
C.2 Dataset Details
C.2.1 Computer Science Dataset.
We first build our scientific research ecosystem using real scientists’ information from the AMiner Computer Science dataset  [^6] https://www.aminer.cn/aminernetwork, which was constructed by extracting scientists’ profiles from online web databases (Tang et al., 2008). This dataset includes 1,712,433 authors and 2,092,356 papers, covering the period from 1948 to 2014, with disambiguated author names. To manage the large volume of data, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2000, 2010, and 2014, respectively. For quality assurance, we filtered out past papers lacking abstracts or with fewer than 10 citations, contemporary papers with fewer than 5 citations or missing abstracts, and authors with fewer than 50 papers or 50 co-authors. As a result, we extracted detailed information from 156 authors and 178,197 papers to construct the ecosystem (85,217 papers for the past database and 92,980 papers for the contemporary database) and initialize the corresponding agents for the simulation. All paper and author data are embedded using the “mxbai-embed-large” model (Lee et al., 2024).
C.2.2 Cross-domain Dataset.
To demonstrate the generalizability and robustness of the system, we conduct additional experiments on the Open Academic Graph 3.1  [^7] https://open.aminer.cn/open/article?id=65bf053091c938e5025a31e2, which developed from the Open Academic Graph (Zhang et al., 2022). This dataset includes 35,774,510 authors and 130,710,733 papers as of 2023, spanning diverse domains such as physics, chemistry, computer science, and biology. To manage this extensive dataset, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2010, 2020, and 2023, respectively. For quality control, we removed past papers without abstracts or with fewer than 200 citations, contemporary papers with fewer than 50 citations or missing abstracts, and authors with fewer than 40 papers or 50 co-authors. This filtering resulted in detailed information on 3,169 authors and 201,131 papers to construct the ecosystem (139,646 papers for the past database and 61,485 papers for the contemporary database) and initialize agents for simulation. Unlike the Computer Science dataset, the Open Academic Graph lacks author-level information on research interests, citations, and co-authors. To address this, we extracted each author’s published papers between $y_{start}$ and $y_{bound}$ and enriched their profiles with the keywords, citations, and authors associated with these papers. Given that the complete set of keywords from all an author’s published papers could be overwhelming, we prompted GPT-4o (OpenAI, 2023) to extract and summarize these keywords into concise research interests for each author. All paper and author data were similarly embedded using the “mxbai-embed-large” model.
C.3 Detailed Comparison Settings
Given the problem formulation of the AI Scientist, HypoGen, and ours, we must make several justifications to ensure relative fairness in our comparisons: (1) Since the AI Scientist is limited to generating ideas from predefined topics (2D Diffusion, NanoGPT, and Grokking), we include NanoGPT in the topic selection prompt for both HypoGen and VirSci as the initial discussion topic, ensuring that the final abstracts align with the same research direction. (2) Given the different approaches to idea generation, we ensure that comparisons are made under similar inference costs. The AI Scientist performs 50 turns of self-reflection during its idea generation, which does not apply to its paper (abstract) generation. To match the inference costs, we set VirSci with 4 team members and 5 discussion turns, and set HypoGen to 12 discussion turns, ensuring that the experiments are conducted under comparable computational costs. (3) Since the AI Scientist and HypoGen lack a scientific research ecosystem, they retrieve papers across all time ranges through the Semantic Scholar API (Fricke, 2018) or PubMed [^8] https://pubmed.ncbi.nlm.nih.gov/. To maintain consistency, we replace our databases and HypoGen’s PubMed API with the Semantic Scholar API for paper retrieval and metric calculation. Specifically, after generating ideas and corresponding abstracts, we use the generated ideas as queries to retrieve related papers, extracting the relevant abstracts and citation counts for evaluation. (4) We evaluate the generated abstracts using both our metrics (CD and CI), the AI Scientist’s metric (LLM review score) (Lu et al., 2024) and the human evaluation (Si et al., 2024). The LLM review score is calculated by GPT-4o, which conducts abstract reviews based on a truncated version of the Neural Information Processing Systems (NeurIPS) conference review guidelines [^9] https://neurips.cc/Conferences/2024/ReviewerGuidelines, shown in Fig. 25.
@@ C.3 Detailed Comparison Settings
C.4 Human Evaluation
To comprehensively compare our method with baseline models, we additionally conduct a human evaluation to assess novelty, feasibility, and effectiveness besides our proposed metrics and LLM review score. Ten PhD students in computer science (the relevant field for the selected research topic), unaware of the method identities, rated the abstracts on a 10-point scale (1: Poor, 10: best) based on three metrics (Si et al., 2024): 1. Novelty: Whether the idea is creative and different from existing works on the topic, and brings fresh insights; 2. Feasibility: How feasible it is to implement and execute this idea as a research project; 3. Effectiveness: How likely the proposed idea is going to work well (e.g., better than existing baselines). The 10-point scale guideline is provided for human experts and is designed based on the principles outlined by Si et al. (2024).
Team Size Invitation Mechanism Turns
@@ C.4 Human Evaluation
Team Size Novelty Assessment Turns
@@
Team Size Discussion Topology Turns
1 2 3 4 5 6 7
4 Sequential Mode 1.75 0.90 2.36 2.00 3.40 2.91 2.53
Random Mode 1.67 0.85 2.34 2.01 3.35 2.84 2.47
8 Sequential Mode 3.48 3.67 3.97 3.56 4.23 3.48 3.43
Random Mode 3.50 3.63 3.88 3.52 4.16 3.40 3.39
Appendix D Ablation Study
D.1 Effects of Paper Database
In the workflow of the proposed VirSci, the past paper database is utilized as a reference during idea generation and novelty assessment. To verify the role of the constructed paper database, we conduct ablation studies on both datasets under the setting of 8 team members and 5 discussion turns, as shown in Tab. 3 and 4. The experimental results demonstrate that having references enhances the novelty of the final generated ideas, avoiding shallow or fanciful outcomes produced by the system.
D.2 Effects of Components Designed for Improving Novelty
D.2.1 Invitation Mechanism
D.2.2 Novelty Assessment
D.2.3 Self-review
@@
Team Size Pattern Turns Cost $\downarrow$ ON $\uparrow$
@@
D.3 Effects of Discussion Pattern
D.3.1 Discussion Topologies
D.3.2 Discussion Turns
In the previous experiments, we fixed the number of discussion turns in each step to ensure fair comparisons. However, in real-world research environments, teams of scientists do not spend the same amount of time on each stage of the research process. To explore this, we compare fixed discussion turns with adaptive turn numbers. In the adaptive pattern, the team leader decides whether the team needs additional turns based on the current progress and the goals of each stage. The results of both patterns, along with their corresponding inference cost, are shown in Tab. 9. The comparison reveals that the adaptive pattern achieves a higher ON while reducing the inference cost. This efficiency can be attributed to the more flexible approach, allowing teams to adjust their discussions dynamically rather than adhering to a rigid structure (which may lead the team in the wrong direction when a section is over-discussed). Furthermore, examining the number of turns at each stage in both 4-person and 8-person teams under the adaptive pattern offers additional insights. Larger teams require more discussion turns and face greater challenges in reaching consensus (Janis, 1972; Pitters and Oberlechner, 2014). This highlights the adaptive pattern’s advantage in accommodating the complexities of larger teams while maintaining a higher level of novelty in the final research outputs.
D.4 Effects of Different Exploration Mechanisms on Scientific Collaboration
If we regard the collaboration among our scientists as an "explore-exploit" model, collaborating with scientists they have not worked with before can be viewed as “explore", while collaborating with previously partnered scientists can be seen as “exploit". In this section, we further investigate how the mechanism of "explore" impacts our final results.
In the default setting, when creating the adjacency matrix, we add 1 to each value in the matrix to ensure that agents collaborate with individuals they have not worked with before, treating "explore" as uniformly distributed. Based on this, we also conducted experiments where the incremental values follow a normal distribution, defined as: